Dublin University Rifle Club

D.U.R.C. Members Handbook

Contents

1 Message from the Captain 1

2 Safety 3

3 Club Rules and Regulations 3 3.1 Safety Regulations . . . . . . . . . . . . . . . 3 3.2 General Regulations . . . . . . . . . . . . . . 5

4 Booking Details 9

5 Free Rifle 10

6 Air Rifle 11

7 Sporter Rifle 12

8 Shooting positions 13 8.1 Prone . . . . . . . . . . . . . . . . . . . . . . 13 8.2 Standing . . . . . . . . . . . . . . . . . . . . . 14 8.3 Kneeling . . . . . . . . . . . . . . . . . . . . . 15

9 Competitions 16

10 Club Competitions 18 10.1 Ladders . . . . . . . . . . . . . . . . . . . . . 18 10.2 Air Rifle Championships . . . . . . . . . . . . 18 10.3 Free Rifle Championships . . . . . . . . . . . 19 10.4 Sporter Rifle Championships . . . . . . . . . . 19 10.5 Rapid-Fire Competition . . . . . . . . . . . . 20 10.6 Turkey Shoot . . . . . . . . . . . . . . . . . . 20 10.7 DURC 10m Air Rifle Open . . . . . . . . . . . 20 10.8 DURC 50m Free Rifle Open . . . . . . . . . . 21 10.9 DURC/UCDRC Air Rifle Colours . . . . . . . . 21

11 Club Awards 23 11.1 Standards . . . . . . . . . . . . . . . . . . . . 23 11.2 Club Colours . . . . . . . . . . . . . . . . . . 23 11.3 Best Newcomer Award . . . . . . . . . . . . . 24

12 Competitions Calendar 2000-2001 25

13 The DURC Squad 27

14 Contact Points 28

15 Who’s Who 29 15.1 Your Committee . . . . . . . . . . . . . . . . . 29 15.2 Your Range Officers . . . . . . . . . . . . . . . 29

16 College Map 31

1 Message from the Captain

1 Message from the Captain

Welcome to the Rifle Club !

Target Shooting has seen a great year in the Irish shooting community at large, with Nicolas Flood from Fassaroe taking the world cup gold medal in Mexican Silouette shooting, and with our Olympic shooters, Alan Lewis (50m Free Rifle Prone), David Malone and Derek Burnett (Men’s Trap, or Clay Pigeon) doing very well in the face of enormous pressure. For DURC, it has been an excellent year as well, with several promising newcomers like Sharon Wall and Aisling Broderick, as well as old hands doing very well in National competitions, where they have dominated their various classes at every competition we have attended.

This year promises to be one of our best yet. The range has had some redecorating work done and looks a lot better for it. We have a full year of competitions planned, both internal and external, and this year sees the DURC Squads restarted after an absence of a few years. The Squads were the core of our competitive shooters and it will be interesting to see the competition this year as the new up-and-coming

The Club Website, which was reworked last year by Mark Dennehy, is looking forward to even more great changes, including discussion boards, webcam broadcasts of Squad training and competitions, and this year we hope to test online booking to make it even more convienent for college students to book details. That should be making an appearance around the start of Hilary Term for an experimental period.

This year sees the first year in some time that UCD Rifle Club has a colours team without the benefit of National Squad members, so the playing field is now level and we look forward to giving them a severe trouncing! The colours is the main UCD-TCD competition and social event of the year - in fact we hold it about twenty paces from the bar to ensure this ! So get training soon and don’t think that just because you aren’t picked to compete that you can’t come - this is one of our biggest parties and all are welcome, so even if you’re not picked, come along for the craic !

Lastly, don’t be afraid to try to get involved with the club - we can always use a few more good people to help out with running DURC. Remember, it’s your club !

Good Shooting,

John Keeney, Club Captain

2 Safety

2 Safety

The club has always made safety its primary goal. As such, it is expected that every member will familarise himself/herself with the club’s safety regulations. No member is permitted to book a detail without having first either attended the safety meeting held at the start of the year, or one of the safety briefings which are held every week throughout the year.

Further, you should note that shooting may look like it is not strenous, and in fact it does not require onerous amounts of physical strength, but you are lifting an eight-kilo rifle into your shoulder every time you fire and stretching muscles that are not normally stretched - so WARM UP FIRST!!!. Walk briskly for a few minutes to get to the rifle range, and once there, stretch carefully to get used to the rifle position. If in doubt as to the best way, ask a Range Officer. You might feel a little silly - but you’ll feel much worse if you strain a back muscle while shooting !!!

3 Club Rules and Regulations

All members are required to acquaint themselves with all regulations. Ignorance of rules will not be accepted as an excuse. If in doubt, ask a Range Officer (R.O.).

3.1 Safety Regulations

1. Never, under any circumstances, point a rifle - whether loaded or not - at anyone.

3 Club Rules and Regulations

2. Except when actually engaged in shooting, the rifle bolt must be kept fully open and the rifle unloaded. Where appropriate, magazines should be removed. Air Rifles should remain uncharged, unloaded and with the action open until shooting.

3. Target changing may be carried out only with the permission of the Range Officer who will ensure, having called a ceasefire - that all actions are open. Where target changers are used for Air Rifles, it is deemed polite to refrain from changing targets if someone at the firing point is shooting.

4. Members will not commence shooting until the Range Officer has given the ”Commence Fire” instruction. They must stop immediately once the ”Cease Fire” instruction is given.

5. Members must ensure that the rifle is pointing downrange at all times.

6. No member may remove a rifle from the firing point, unless directed to do so by the Range Officer.

7. In the event of any problem arising with a rifle, e.g. a round failing to detonate, call the Range Officer. Never attempt to deal with the problem yourself. He/She has been trained to handle it safely.

8. Rifles must be carried with the bolts or actions visibly open and in a safe manner. E.g. by the center of gravity of the rifle and not slung over a shoulder.

9. .22 rounds must be kept in the blocks provided and used only as required. These rounds are rimfire and so may detonate if dropped on a hard surface.

3.2 General Regulations

10. Smoking on the range or near ammunition is prohibited.

11. The consumption of alcohol before shooting is expressly prohibited. Any member who arrives at the range expecting to shoot is required to be free from the influence of alcohol or any other controlled substance, and will not be permitted to enter the range unless free from said influences.

12. Only Range Officers and those actually engaged in shooting may be present on the firing point during a detail. No one shooting on a detail may leave the firing point without permission from a Range Officer.

13. Members are required to familiarise themselves with the Club’s General Regulations.

3.2 General Regulations

1. All members, before they are allowed to shoot, must

• Have joined previous to that detail,

• Have attended either the Safety Meeting held at the start of Michaelmas term, or one of the weekly Safety Briefings, and have satisfied the Instructor that they are familiar with the Club’s safety regulations.

2. Booking of details :

• Details are booked on the grids provided at our front gate noticeboard. Each detail is of 30 minutes duration.

• You may not book more than one detail per week in Michaelmas term unless otherwise informed by the Club.

• Unless you are otherwise informed by a Range Officer, you may not book more than one detail per day.

• Squad training sessions are not counted as details, and so a squad member may attend squad training and book a detail in the same week, even in Michaelmas term.

• Members may, if they expect details to fall vacant, turn up on the range in the hope of getting a free detail. Whether he/she does will depend on the R.O., who will advise in accordance with the Committee’s directions.

• Members are required to turn up at the range at least 15 minutes before their detail is due to start. Those who arrive late for their detail forfeit their shooting time. Later details will not be delayed to account for lateness.

• Should a member be more than 15 minutes late past the start of his or her detail, that detail will be vacated. Should the member be more than 30 minutes late, it will be noted as a no-show.

• If two consecutive details fall vacant, the R.O. may terminate shooting for the night.

• Unless you are sure that you can turn up, do not book a detail.

• Should you have booked a detail and you subsequently find that you cannot attend, delete your name from the grid as soon as possible, thus enabling someone else to take your place. Failure to turn up for a detail which you have booked will lead to your shooting privileges being revoked until such time as a suitable apology is offered. You may not book details until after this apology is offered.

3. Wilful damage to club property is a serious offence and will be treated as such.

4. Most items of Club equipment, especially the rifles, are precision instruments. Members must ensure that they treat the equipment with the care and respect it deserves : e.g. don’t use excessive force in opening and closing of rifle bolts.

5. Unless otherwise directed, members should replace equipment in it’s proper place after use. This includes washing of mugs after coffee. This does not include removing rifles from the firing point after your detail unless directed to do so by the Range Officer.

6. Each member is expected to clean up his/her shells and change his/her targets after the detail.

7. Club files are confidential and may be consulted by Range Officers or Committee Officers only. Should you wish to see your records, you are entitled to do so, but a Range Officer or Committee member must retrieve them for you.

8. Scoring is the job of the Range Officer / Statistics Officer and will be carried out in accordance with I.S.S.F. /

N.R.P.A.I. / N.T.S.A. rules.

9. Ammunition will be issued through and recorded by the Range Officer on duty only.

10. Members are expected to pay for ammunition each night at current rates.

11. Members must comply with all reglations and all lawful directions or instructions of the Range Officer(s) while on the range.

12. Access to the range, whether by a member or nonmember, is at the discretion of the Range Officer, who may withdraw that permission at any time subject to Article 15B of the Club constitution.

13. Actions deemed by the Range Officer(s) to be in breach of the Safety Regulations or General Regulations may be punished by suspension or expulsion from the Club and/or by prosecution as outlined in Articles 15C, 19A and 19B of the Club Constitution.

14. Requests for letters or recommendations re firearms certificates will only be given consideration by the Committee in cases of the applicant being an active member of at least one year’s standing. The Committee reserves the right of refusal.

15. In addition to the D.U.R.C. Club regulations, those contained in the Firearms Acts currently in force also apply to members.

16. The omission of regulations or directives from these regulations (Safety and General), does not mean that

4 Booking Details

others cannot be enforced. The Committee reserves the right to add to or alter the regulations of this Club.

17. Members who have suggestions / queries may voice them by :

• discussing the matter with a Range Officer

• contacting the Committee

• posting their suggestions or queries on the general email mailing list maintained by the Club

D.U.R.C. Committee, 2000

4 Booking Details

Due to the number of members in the club (usually between three and five hundred), you cannot simply arrive at the range and hope to shoot for an hour or so. Time on the range is divided into details of a half-hour, beginning at 1800 on weekdays during term time. The last detail is either the last one booked that night or the one beginning at 2130, whichever is earliest.

Due to the demand for time on the range, there are some rules for booking details and these are laid down in the Club’s General Regulations above.

5 Free Rifle

5 Free Rifle

In smallbore (.22 calibre) shooting, competitors shoot from prone, standing and kneeling positions at a firing line 50 metres from the target. There is also a prone-only competition, which is what we mainly shoot due to constraints of space, however it is perfectly possible to train for both. Free Rifle referes to the class of rifle used. For Standard Rifle competitions, you may not modify the rifle (hence it is a ’standard’ rifle for all competitors), whereas you are free to do so for Free Rifle. Standard rifle competitions are some-

6 Air Rifle

what rare in Ireland, though Sporter Rifle is somewhat similar. The bullseye is just 10.4 mm in diameter and the Olympic record for the prone rifle event is a maximum 60 bullseyes with 60 shots. In DURC, we don’t have a 50m range and so shoot the indoor variant at 25 yards with correspondingly smaller targets. This is by no means a ’second-best’ arrangement, in fact 25 yard indoor competition is a great deal more exacting, and competitors who may shoot perfect scores regularly in 50m competitions are hard-pressed to shoot perfect 25 yard cards. We do train for, and compete successfully in 50m competitions, but at other ranges.

6 Air Rifle

Air Rifle shooters fire .177 calibre (4.5mm) pellets from the standing position using compressed air rifles. The air rifles we use are not completely dissimilar from ”BB guns”, but are more powerful and accurate, not to mention a lot more expensive ! The targets for Air Rifle are fired at from the standing position at a range of 10 meters. The bullseye is 0.5mm wide - that’s about as large as the full stop at the end of this line. The World Record is currently 600 - that’s 60 perfect shots in a row !

7 Sporter Rifle

7 Sporter Rifle

Sporter Rifles are mainly used in Mexican Silouette Shooting outdoors. This involves shooting at metal silouettes at varying ranges. This year we are hoping to enter a squad at the national level competitions that take place around the country, which will be the first time in some years that we have done so. The actual competitions are shot from the standing position at ranges up to 300 metres, using telescopic sights. The rifle is not customisable, and is the only branch of target shooting that we currently do where that is the case. This means that you must now tune yourself instead of your rifle !

8 Shooting positions

8 Shooting positions

8.1 Prone

The prone position is the main shooting position which we train for with Free Rifles in DURC. The basics of the position are simple : lie down at an angle to the target, with your legs spread for stability and use the sling to take the bulk of the weight of the rifle.

While simple to start with however, tuning your position to allow you to get that final few points takes time, patience and experimentation.

8.2 Standing

Standing (or Offhand), is the hardest position to master. It is the only position we train in for Air Rifle. The basics are to stand with your feet a comfortable distance apart, your side facing the target. The rifle is supported by the left arm if you shoot righthanded, and it rests on the left hip. This is where most women have a distinct advantage in that men’s hips tend to be somewhat less useful as an elbowrest ! The objective is to have a solid line of bone from the hand to the toes supporting the rifle.

It will take time to master this position, as your muscles will have to adapt to the unique stresses it presents. However, Air Rifle is the event that the Colours take place in, and for the more ambitious, Nancy Johnson was able to take Olympic Gold in this event this year with only two years of competition practise !

8.3 Kneeling

8.3 Kneeling

Kneeling is the third position in Three-Position shooting. Competitors kneel, using a padded kneeling roll under their ankle, and support the rifle on their arm, which rests on their upright knee. The sling may also be used. This position tends to stress the knees and ankles - be sure to always warm up before attempting it !

9 Competitions

9 Competitions

Each competition will have its own set of rules, whether ISSF, NTSA or NSRA or those decided upon by some other governing body. The minutae of these are too detailed for this handbook and for 95% of the time do not even arise. The basic rules are relatively straightforward and center on the following concepts:

• Safety. At all times, in all competitions, safety is the primary rule. At all times listen to the Range Officer in charge and obey his instructions to the letter. Do not go forward of the firing line without his express permission. If you have any doubt as to what to do, ask the Range Officer first !

• Good Sportsmanship. This should go without saying, but we’re going to say it anyway. In competitions you are not just representing yourself, but also your Club and possibly your Country. As such, exhibiting good sportsmanship is a necessity as well as being polite. Being a good winner is as important as a good loser - remember that if you win the competition it only means that the person who could beat you wasn’t there that day !

• Self-sufficency. A general rule is that leaving the firing point in the middle of the detail is prohibited. So before you settle down to the task of shooting, check to see that you have everything you will need - especially enough ammunition !

• Silence. Yelling on a firing range is decidedly a bad idea, somewhat akin to yelling ’Fire’ in a crowded theatre. It distracts the other competitors, usually results in the Range Officer being convinced that a serious accident has happened and is never considered funny. When you are shooting, remain quiet. (This also applies to muttering to yourself about the quality of your shooting)

• Focus. Even if you are just out to enjoy yourself, which is perfectly approporiate and the main goal of all involved, many shooters tend to try to bring themselves to focus on their shooting before a match. So if someone is about to go to their firing detail, don’t disturb them unless you’ve allready been chatting away. The National Squad in particular tend to train quite seriously, and while they enjoy the craic as well as everyone else, a few prefer to have a few minutes before they start to get focussed.

• Fun. It’s not often pointed out, but at the end of the day this is not a college academic course - it’s a sport or hobby. It’s meant to be fun. At competitions, enjoy the social aspect as well as learning from the other shooters. Enjoy the after-match socialising, in fact enjoy the whole sport - the training, the preperation, the competition and the parties too !

10 Club Competitions

10 Club Competitions

All internal club competitions are held on-campus in the DURC range and include two classes - Experienced for those who have shot in that discipline for more than one year, and Novice for those who are on their first year. Note that you may be Experienced in one discipline and a Novice in another ! All prizes and results are officially presented at the AGM in Trinity Term, which is usually followed by a few celebratory drinks!

10.1 Ladders

From your third or fourth detail (your Range Officer will say when), you must shoot a competition card for the internal Ladders at each detail. This comprises one ten-target card for Free Rifle, a five-string card with two rounds per target for Air Rifle, or three cards of 5 rounds each (one card per position) for Sporter Rifle. These scores are tracked throughout the year and will be listed on the Club’s noticeboards in the range and also on the Website. Prizes will be awarded to the top three shooters in each ladder at the AGM and consist of shields for the top three experienced class shooters and medals for the top three novice class shooters. The Ladders are also the means by which the Squads are selected so do your best at them !

10.2 Air Rifle Championships

The Air Rifle Championships are held in Trinity Term. They consist of up to five sighters and twenty competition shots to be fired in the course of a 30 minute detail. The overall

10.3 Free Rifle Championships

score is out of 200 and last year’s experienced class winner was Ray Reilly with 174, and last year’s novice class winner was Sharon Wall with an exceptional 152. Medals will be presented to the top three shooters in both classes at the AGM.

10.3 Free Rifle Championships

The Free Rifle Championships are held in Trinity Term. They consist of five sighters and twenty competition shots to be fired in the course of a 30 minute detail. The overall score is out of 200 and last year’s winners were John Keeney in Experienced class with 188 (a score shared with Ian Marnane but John won on countback, which shows how close that competition was !), and Anna Hickey in Novice class with a spectacular 186 ! Again, medals will be presented to the top three shooters in both classes at the AGM.

10.4 Sporter Rifle Championships

The Sporter Rifle Championships are held in Trinity Term, and consist of five sighters and fifteen competition shots (five in each of the three positions), all in 30 minutes. The overall score is out of 150 and last year’s experienced class winner was Willie O’Connor with 137. No entries for the Novice class arrived last year, so this year give it a try - it might prove to be an easy medal ! Again, as with all competitions, the medals get awarded to the top three shooters in each class at the AGM.

10.5 Rapid-Fire Competition

The rapid-fire competition is most definitely not an international standard competition, but has proved over the years to be the most enjoyable competition for most of the casual DURC members. Held in Trinity Term, it is the last DURC competition of the academic year. You have up to five minutes to use your five sighters, and then you have to complete a ten-target card in ninty seconds ! That works out as ten seconds per shot as the time starts from the first fired shot. However, this year sees a rule change to make things more interesting : any time remaining at the end of the ninty seconds gets added to your score, with every two seconds being worth an additional point. However, tiebreaks are settled by comparison of the actual card score, so don’t be too reckless ! Again, medals are awarded to the top three shooters in each class at the AGM.

10.6 Turkey Shoot

The Turkey shoot is a traditional christmas novelty competition where silouettes of chickens and other targets are shot at with air rifles. Mulled wine to be consumed after the shooting ! A Plaque is awarded to the top shooter in each class, with medals for the second and third places.

10.7 DURC 10m Air Rifle Open

The 10m Air Rifle Open is the annual open competition where DURC members get to see how well they match up to the Irish shooting community at large. This year we’re holding the open in UCD due to technical difficulties, but there will

10.8 DURC 50m Free Rifle Open

be white chocolate cheesecake and lots of prizes for various classes to compensate !

10.8 DURC 50m Free Rifle Open

The Free Rifle Open is the Free Rifle version of the Air Open. Traditionally the first competition of the 50m season, it is shot in Fassaroe Shooting Club and along with the many prizes, there is a barbecue held concurrently so that hungry shooters coming off the firing line can tuck into a homecooked hamburger or chicken kebab ! (And of course, there are the odd few vegetarian burgers as well)

Note that both opens are registered shoots - so the scores you get there could qualify you for selection to the National Squad and international competitions if you are good enough !

10.9 DURC/UCDRC Air Rifle Colours

And lastly, but by no means least, there is the Colours match. Forget the rest, this is where the honour of trinity is at stake ! DURC versus UCD in the Wilkinstown shooting club, and the fact that Wilkinstown shooting club is twenty paces from a pub is convienent for the celebrations after we win, and for Geoff’s famous soup to warm shooters up beforehand ! Usually this is a serious social event with the National Squad and others showing up for the celebrations, and most of DURC tend to stay overnight and return the following morning. Transport to and from Wilkinstown is by a bus shared by UCD and DURC so the pre-match joshing and prediction can begin good and early. Last year saw DURC’s novice team triumphant, but a sneaky move by UCD (recruiting two of the National Squad for their team) meant that our Experienced team lost by a narrow margin. To compete in the colours this year, you have to be good enough, so get training now ! And if the thought of representing DURC against the ancient enemy (smile, this is supposed to be a fun competition !) is not enough to stir the blood, every colours competitor recieves a commemurative medal at the AGM !

11 Club Awards

11 Club Awards

Along with the Club competitions, there are the Club awards, which are presented for outstanding performance throughout the year, or for exceptional skill as in the case of the Standards.

11.1 Standards

The Standards are awarded for achieving above a certain standard of skill, demonstrated by shooting above a set score in three consecutive target cards inside of one hour. There are five badges given out to denote the different standards :

• StarShot

• TopShot

• SharpShot

• CrackShot

• HotShot

• SureShot

To find out more, ask the Match Secretary.

11.2 Club Colours

Club Colours are awarded at the AGM, at the captain’s discretion, to a member who has, in the captain’s judgement, performed to a remarkably high standard. Last year saw Sharon Wall awarded Colours for the quite astonishing feat of going from complete novice to winning Class C in the Air Rifle National Championships in under eight months, which is almost unheard-of.

So train seriously, and you might be getting colours this year !

11.3 Best Newcomer Award

This is awarded to the best novice of the year, according to Committee decision. Last year saw Anna Hickey receiving the award, an engraved perpetual silver tray, for her exceptional improvement in Free Rifle, where she won the novice section of the internal championships with a score only two points off the winner of the experienced class ! So if you’re only joining the club this year for the first time, enter as many competitions as you can - you might have your name engraved as the best newcomer of the year 2000 !

12 Competitions Calendar 2000-2001

12 Competitions Calendar 2000-2001

Note that this calendar may be changed during the year. If this happens, notices will be posted on the Email List, the Noticeboard in front arch, the noticeboards in the range and on the Website.

Date  Event  Venue

October 2000

2 – 6  Fresher’s Week
9  Michaelmas Term starts
9 Safety and Orientation Meeting Davis Theatre
14, 15 10m Wilkinstown Open 1  Wilkinstown
16  DURC Shooting Begins
28,29  10m UCD Open 1  UCD

November 2000

3  Initial Squads selection announced
5  50m Leinster Open  Blessington
11,12  10m WAIDAM (Fassaroe Shooting Club) Open  TBA
26  10m DURC 10m Air Rifle Open  UCD

December 2000

6  DURC Turkey Shoot  TCD
8  Michaelmas Term ends

January 2001

8  Hilary Term starts
13,14  10m Wilkinstown Open 2  Wilkinstown
27,28  10m Fassaroe Shooting Club Air Open  TBA

12 Competitions Calendar 2000-2001

February 2001

11  25yd  Rathdrum Rifle Club Open  Rathdrum
17,18  10m  UCD Open 2  UCD

March 2001

3  10m Air Rifle Colours  Wilkinstown
9  Hilary Term ends
10,11  10m  Air Rifle Nationals  UCD
25  25yd  Rathdrum Rifle Club Spring Special  Rathdrum

April 2001

7  50m  NTSA Three-Position Event  FSC, Enniskerry
8  50m  DURC Free Rifle Open  FSC, Enniskerry
9  Trinity Term starts
11  25yd  DURC Sporter Rifle Championships  TCD
15  25yd  Indoor Free Rifle Nationals  Rathdrum
18  10m  DURC Air Rifle Championships  TCD
21  10m 50m Isle of Man Shoot  Isle of Man
25  25yd  DURC Free Rifle Championships  TCD
29  50m  Fermoy Open  Fermoy

May 2001

2  25yd  DURC Rapid-Fire Competition  TCD
6  50m  Co.Wicklow Open  Blessington
9  DURC AGM  TBA
11  Trinity Term ends

13 The DURC Squad

13 The DURC Squad

The DURC Squads are the best competitors we can field to external competitions. Squad members get to train for an extra detail or so per week during the squad training times (and are in fact expected to attend these at a satisfactory rate) and are the pool that the colours team is picked from. The Squads comprise the top twelve shooters in each of the three disciplines (Air, Free and Sporter Rifle), but there is also a minimum qualifying score for entry. Scores are taken from the Ladders to select the members of the squad, and if you should consistently out-perform a member of a squad on the ladders, you will be invited to fill his slot on the squad.

And for the novices, don’t worry - the Novice Squads work the same way (though there are only six on the novice squads), and the Novice team for the Air Rifle Colours are selected from the Novice Squads as well.

Men and women compete equally to enter the squads, and age is not counted either. So - do you want to be listed as the best ?

14 Contact Points

14 Contact Points

To contact the club, there are several contact points : • By phone : (01) 608 1281

• By mail : Dublin University Rifle Club, c/o DUCAC, Trinity College Dublin, Dublin 2.

• By email : Rifle.Club@tcd.ie

• By email mailing list : Mail majordomo@tcd.ie and put subscribe durc-list in the body of your message.

• The Website is at http://www.tcd.ie/clubs/rifle

• And our noticeboard is in Front Arch, as noted on the map.

15 Who’s Who

15 Who’s Who

15.1 Your Committee

These are the members who run the club from day to day and do all the logistics, publicising, and all the little menial jobs that keep the club going.

• John Keeney, Captain

• Alan Smyth, Secretary

• Ben Jones, Treasurer

• Willie O’Connor, Chief Range Officer

• Mark Dennehy

• Eric Farrell

15.2 Your Range Officers

These are the Range Officers who supervise shooting each night on the range.

• Willie O’Connor, Chief Range Officer

• John Keeney

• Alan Smyth

• Mark Dennehy

• Eric Farrell

• Keith White • Sharon Wall

• Aishling Broderick • Mark Treanor

• Ray Reilly

• Grainne McDermott

16 College Map

16 College Map Notes